By using the method of undetermined coefficients, find the general solution of the following differential equation

(f) / + 4y = cos 2x.
Answer #1

?) +4a = Cos 21 1. let y= einx 2 m2 + 4 = 0 - м = +25 №3 пе () = е свъ (2) +C25in (2х).Noco yp in of the form Ax Cos (2x) + Bx sin (2x) So, yp = Ax cos(2x) + B x sin(2x) yo= A cos (28) – 24x sin (2x) + Bain (2x)comparing both sides we have 48=i =B = . and 4A =0 » Azo There fone Mp = 0.x cos (2x) + sin(2x) = Lin (2x) Therefore the ge